Question 7 Suppose vx-2t2 and vy-12 t + sin(t) + cos(t) + et. If the initial position at t-0 is x-0, y-o, what is the final x position after 2 seconds? Select the correct answer CHECK ANSWER 0 of 1 attempts used LAST ATTEMPT O Between 0 and2 O Between 8 and 10 O Between 4 and 6 O Between 6 and O Between 2 and 4
Answer #1

vx = 2t2

x = \int 2t2 dt

x = 2t3/3 + C

at t = 0 , x = 0

so,

C = 0

x = 2*t3/3

at t = 2 s

x = 2*(2)3/3

x = 5.33 m

option 3 is right .